OpenStudy (anonymous):

2 + 1 + 1/2 + 1/4 + 1/8+ ... A. 2.25 B. 3.5 C. 4 D. 5 E. infinite

OpenStudy (anonymous):

just add then all together, do it two at a time to make it easier 2+1= 1/2+1/4= 1/8 + the other 2 =

OpenStudy (anonymous):

c.4